# Cybersecurity of Quantum Key Distribution Implementations

### v2

@@ Appendix A Application of Linear Optical Devices on Photonic States
When a single mode, carrying one or more photons, enters the interferometer, three ancillas in a vacuum state are added by the interferometric setup (see Figure 4). As mentioned above, the mode that enters the interferometer at time $t^{\prime}_{0}$, yields two modes at time $t_{0}$, and two modes at time $t_{1}$. These four output modes are: times $t_{0}$, $t_{1}$ at the ‘s’ (straight) arm of the interferometer, and times $t_{0}$, $t_{1}$ at the ‘d’ (down) arm of the interferometer. A basis state of this Fock-space can be written as $|{n_{s_{0}},n_{s_{1}},n_{d_{0}},n_{d_{1}}}\rangle^{\!\text{\tiny F}}$.
(1)(1’) Pulse (1) is about to enter the interferometer. A vacuum ancilla (1’) is added at the input of the first beam splitter, $BS_{1}$.
(3)(2)(2’) Pulses (1) and (1’) interfere in the first beam splitter ($BS_{1}$) and yield a superposition of (2) and (3) in the short and long arms of the interferometer, respectively, $|{1}\rangle^{\!\text{\tiny F}}_{1}|{0}\rangle^{\!\text{\tiny F}}_{1^{\prime}}\stackrel{{\scriptstyle BS_{1}}}{{\longrightarrow}}(|{1}\rangle^{\!\text{\tiny F}}_{2}|{0}\rangle^{\!\text{\tiny F}}_{3}+i|{0}\rangle^{\!\text{\tiny F}}_{2}|{1}\rangle^{\!\text{\tiny F}}_{3})/\sqrt{2}$. Pulse (2) is about to enter the second beam splitter ($BS_{2}$) so a vacuum ancilla is added (2’).
(3’)(3)(4)(5) Pulses (4) and (5) are created by pulses (2) and (2’), $\frac{1}{\sqrt{2}}|{0}\rangle^{\!\text{\tiny F}}_{2^{\prime}}|{1}\rangle^{\!\text{\tiny F}}_{2}\stackrel{{\scriptstyle BS_{2}}}{{\longrightarrow}}(i|{1}\rangle^{\!\text{\tiny F}}_{4}|{0}\rangle^{\!\text{\tiny F}}_{5}+|{0}\rangle^{\!\text{\tiny F}}_{4}|{1}\rangle^{\!\text{\tiny F}}_{5})/2$. Pulse (3) is about to enter the second beam splitter, so a vacuum ancilla is added (3’).
(7)(4)(6)(5) Pulses (6) and (7) are created by the interference of (3) and (3’). $\frac{i}{\sqrt{2}}|{1}\rangle^{\!\text{\tiny F}}_{3}|{0}\rangle^{\!\text{\tiny F}}_{3^{\prime}}\stackrel{{\scriptstyle BS_{2}}}{{\longrightarrow}}(i|{1}\rangle^{\!\text{\tiny F}}_{6}|{0}\rangle^{\!\text{\tiny F}}_{7}-|{0}\rangle^{\!\text{\tiny F}}_{6}|{1}\rangle^{\!\text{\tiny F}}_{7})/2$.
Figure 4: Evolution in time of a single photon pulse through the interferometer with $\phi=0$, $|{1000}\rangle^{\!\text{\tiny F}}_{1,1^{\prime},2^{\prime},3^{\prime}}\to\frac{1}{2}\left(|{1000}\rangle^{\!\text{\tiny F}}-|{0100}\rangle^{\!\text{\tiny F}}+i|{0010}\rangle^{\!\text{\tiny F}}+i|{0001}\rangle^{\!\text{\tiny F}}\right)_{5,7,4,6}$. The output state is denoted by modes $|{n_{s_{0}},n_{s_{1}},n_{d_{0}},n_{d_{1}}}\rangle^{\!\text{\tiny F}}$ that correspond to modes (5), (7), (4), and (6), respectively.
@@
$|{1}\rangle^{\!\text{\tiny F}}_{t^{\prime}_{0}}|{000}\rangle^{\!\text{\tiny F}}\mapsto(|{1000}\rangle^{\!\text{\tiny F}}-e^{i\phi}|{0100}\rangle^{\!\text{\tiny F}}+i|{0010}\rangle^{\!\text{\tiny F}}+ie^{i\phi}|{0001}\rangle^{\!\text{\tiny F}})\thickspace/2\ .$ (14)
Note the three vacuum ancillas that were added. Also, note that a pulse which is sent at a different time (say, $t^{\prime}_{1}$, or $t^{\prime}_{-1}$, etc.) results in the same output state, with appropriate delays. That is, a pulse entering the interferometer at time $t^{\prime}_{i}$ results in the state $(|{1000}\rangle^{\!\text{\tiny F}}-e^{i\phi}|{0100}\rangle^{\!\text{\tiny F}}+i|{0010}\rangle^{\!\text{\tiny F}}+ie^{i\phi}|{0001}\rangle^{\!\text{\tiny F}})\thickspace/2$ in a Fock-space with basis states $|{n_{s_{i}},n_{s_{i+1}},n_{d_{i}},n_{d_{i+1}}}\rangle^{\!\text{\tiny F}}$.
Using the more standard notation of $\{|{t^{\prime}_{i}}\rangle\}$ input states and $\{|{s_{i}}\rangle,|{d_{i}}\rangle\}$ output states used in the rest of the paper, Eq. (14) is given by
$|{t^{\prime}_{i}}\rangle\mapsto\left(|{s_{i}}\rangle+i|{d_{i}}\rangle-e^{i\phi}|{s_{i+1}}\rangle+ie^{i\phi}|{d_{i+1}}\rangle\right)/2.$ (15)
@@
We are now ready to consider the setup of Figure 6 and two input modes, $t^{\prime}_{0}$ and $t^{\prime}_{1}$, that enter the interferometer one after the other, with exactly the same time difference $\Delta T$ as the interferometer’s arms. As a result of this precise timing, the two modes are transformed into a superposition of only six modes (instead of eight modes) at the outputs (see Figure 5). Four (vacuum state) ancillas are added during the process and the resulting six modes are $t_{0}$, $t_{1}$, $t_{2}$ at the ‘s’ arm and the ‘d’ arm of the interferometer. A basis state of this Fock-space is, therefore, $|{n_{s_{0}},n_{s_{1}},n_{s_{2}},n_{d_{0}},n_{d_{1}},n_{d_{2}}}\rangle^{\!\text{\tiny F}}$. If exactly one photon enters the interferometer, we can use Eq. (14) to obtain
$|{1}\rangle^{\!\text{\tiny F}}_{t^{\prime}_{0}}|{0}\rangle^{\!\text{\tiny F}}_{t^{\prime}_{1}}|{0000}\rangle^{\!\text{\tiny F}}$ $\mapsto(|{100000}\rangle^{\!\text{\tiny F}}-e^{i\phi}|{010000}\rangle^{\!\text{\tiny F}}+i|{000100}\rangle^{\!\text{\tiny F}}+ie^{i\phi}|{000010}\rangle^{\!\text{\tiny F}})\thickspace/2$
$|{0}\rangle^{\!\text{\tiny F}}_{t^{\prime}_{0}}|{1}\rangle^{\!\text{\tiny F}}_{t^{\prime}_{1}}|{0000}\rangle^{\!\text{\tiny F}}$ $\mapsto(|{010000}\rangle^{\!\text{\tiny F}}-e^{i\phi}|{001000}\rangle^{\!\text{\tiny F}}+i|{000010}\rangle^{\!\text{\tiny F}}+ie^{i\phi}|{000001}\rangle^{\!\text{\tiny F}})\thickspace/2$ (16)
Recall that $|{0}\rangle=|{10}\rangle^{\!\text{\tiny F}}_{t^{\prime}_{0}t^{\prime}_{1}}$ and $|{1}\rangle=|{01}\rangle^{\!\text{\tiny F}}_{t^{\prime}_{0}t^{\prime}_{1}}$. It follows that an arbitrary qubit is transformed as
@@
(1)(1’)(2) A general single-photon qubit, $\alpha|{10}\rangle^{\!\text{\tiny F}}+\beta|{01}\rangle^{\!\text{\tiny F}}$, enters the interferometer (modes (2) and (1)). Bob adds a vacuum ancilla (1’) that interferes with mode (1) at the first beam splitter ($BS_{1}$).
@@
(2)(2’)(4)(3)(3’) Pulses (1) and (1’) interfere and yield pulses (3) and (4) in the short arm and the long arm respectively, $\alpha|{1}\rangle^{\!\text{\tiny F}}_{1}|{0}\rangle^{\!\text{\tiny F}}_{1^{\prime}}\stackrel{{\scriptstyle BS_{1}}}{{\longrightarrow}}\frac{\alpha}{\sqrt{2}}(|{1}\rangle^{\!\text{\tiny F}}_{3}|{0}\rangle^{\!\text{\tiny F}}_{4}+i|{0}\rangle^{\!\text{\tiny F}}_{3}|{1}\rangle^{\!\text{\tiny F}}_{4})$. Pulse (3) is about to enter $BS_{2}$, so a vacuum ancilla (3’) is added. Pulse (2) is about to enter $BS_{1}$ so a vacuum ancilla (2’) is added.
@@
(6)(5)(4)(7)(8) Pulses (7) and (8) are created by the interference of (3) and (3’) $\frac{\alpha}{\sqrt{2}}|{0}\rangle^{\!\text{\tiny F}}_{3^{\prime}}|{1}\rangle^{\!\text{\tiny F}}_{3}\stackrel{{\scriptstyle BS_{2}}}{{\longrightarrow}}\frac{i\alpha}{2}|{1}\rangle^{\!\text{\tiny F}}_{7}|{0}\rangle^{\!\text{\tiny F}}_{8}+\frac{\alpha}{2}|{0}\rangle^{\!\text{\tiny F}}_{7}|{1}\rangle^{\!\text{\tiny F}}_{8}$. Pulses (5) and (6) are created by the interference of (2) and (2’) in $BS_{1}$ $\beta|{1}\rangle^{\!\text{\tiny F}}_{2}|{0}\rangle^{\!\text{\tiny F}}_{2^{\prime}}\stackrel{{\scriptstyle BS_{1}}}{{\longrightarrow}}\frac{\beta}{\sqrt{2}}(|{1}\rangle^{\!\text{\tiny F}}_{5}|{0}\rangle^{\!\text{\tiny F}}_{6}+i|{0}\rangle^{\!\text{\tiny F}}_{5}|{1}\rangle^{\!\text{\tiny F}}_{6})$.
@@
(6’)(6)(10)(7)(9)(8) Pulses (9) and (10) are created by the interference of (4) and (5) in the second beam splitter $\frac{i\alpha}{\sqrt{2}}|{1}\rangle^{\!\text{\tiny F}}_{4}|{0}\rangle^{\!\text{\tiny F}}_{5}+\frac{\beta}{\sqrt{2}}|{0}\rangle^{\!\text{\tiny F}}_{4}|{1}\rangle^{\!\text{\tiny F}}_{5}\stackrel{{\scriptstyle BS_{2}}}{{\longrightarrow}}\frac{i(\alpha+\beta)}{2}|{1}\rangle^{\!\text{\tiny F}}_{9}|{0}\rangle^{\!\text{\tiny F}}_{10}+\frac{\beta-\alpha}{2}|{0}\rangle^{\!\text{\tiny F}}_{9}|{1}\rangle^{\!\text{\tiny F}}_{10}$ . Pulse (6) is about to enter $BS_{2}$, so a vacuum ancilla is added (6’).
@@
(12)(11)(10)(7)(9)(8) Pulses (11) and (12) are created by the interference of (6) and (6’) in $BS_{2}$ $\frac{i\beta}{\sqrt{2}}|{1}\rangle^{\!\text{\tiny F}}_{6}|{0}\rangle^{\!\text{\tiny F}}_{6^{\prime}}\stackrel{{\scriptstyle BS_{2}}}{{\longrightarrow}}\frac{i\beta}{2}|{1}\rangle^{\!\text{\tiny F}}_{11}|{0}\rangle^{\!\text{\tiny F}}_{12}-\frac{\beta}{2}|{0}\rangle^{\!\text{\tiny F}}_{11}|{1}\rangle^{\!\text{\tiny F}}_{12}$.
@@
Figure 5: Evolution in time of two modes through the interferometer with $\phi=0$, $\left(\alpha|{1}\rangle^{\!\text{\tiny F}}_{1}|{0}\rangle^{\!\text{\tiny F}}_{2}+\beta|{0}\rangle^{\!\text{\tiny F}}_{1}|{1}\rangle^{\!\text{\tiny F}}_{2}\right)|{0000}\rangle^{\!\text{\tiny F}}_{1^{\prime},2^{\prime},3^{\prime},6^{\prime}}\to\bigl{(}\frac{\alpha}{2}|{100000}\rangle^{\!\text{\tiny F}}+\frac{\beta-\alpha}{2}|{010000}\rangle^{\!\text{\tiny F}}-\frac{\beta}{2}|{001000}\rangle^{\!\text{\tiny F}}+\frac{i\alpha}{2}|{000100}\rangle^{\!\text{\tiny F}}+\frac{i(\alpha+\beta)}{2}|{000010}\rangle^{\!\text{\tiny F}}+\frac{i\beta}{2}|{000001}\rangle^{\!\text{\tiny F}}\bigr{)}_{8,10,12,7,9,11}$. The output state is denoted by modes $|{n_{s_{0}},n_{s_{1}},n_{s_{2}},n_{d_{0}},n_{d_{1}},n_{d_{2}}}\rangle^{\!\text{\tiny F}}$.
@@
While a pulse can arrive at the interferometer at any instance of time (which is continuous), we know that pulses only interfere with each other if their time difference is equal to the time difference of the interferometer’s two paths. Thus, we define some instance of time as $t=0$, and consider all times $\{t\cdot\Delta\ \;|\;t\in\mathbb{Z}\}$, where $\Delta$ is the interferometer’s time difference.
@@
$\operatorname{span}\{|{a_{t}}\rangle,|{b_{t}}\rangle\;|\;t\in\mathbb{Z}\},$ (18)
@@
$\operatorname{span}\{|{s_{t}}\rangle,|{d_{t}}\rangle\;|\;t\in\mathbb{Z}\},$ (19)
while we have already shown that the evolution for $|{a_{t}}\rangle$ is given by:
$U|{a_{t}}\rangle=\left(|{s_{t}}\rangle+i|{d_{t}}\rangle-e^{i\phi}|{s_{t+1}}\rangle+ie^{i\phi}|{d_{t+1}}\rangle\right)/\thickspace 2.$ (20)
@@
$U=\frac{1}{2}\begin{pmatrix}1&i&0&0&\cdots\\ i&-1&0&0\\ -e^{i\phi}&ie^{i\phi}&1&i\\ ie^{i\phi}&e^{i\phi}&i&-1\\ 0&0&-e^{i\phi}&ie^{i\phi}\\ 0&0&ie^{i\phi}&e^{i\phi}\\ \vdots&&&\end{pmatrix}$. (22)
@@
$U^{-1}=U^{\dagger}=\frac{1}{2}\begin{pmatrix}1&-i&-e^{-i\phi}&-ie^{-i\phi}&0&0&\cdots\\ -i&-1&-ie^{-i\phi}&e^{-i\phi}&0&0\\ 0&0&1&-i&-e^{-i\phi}&-ie^{-i\phi}\\ 0&0&-i&-1&-ie^{-i\phi}&e^{-i\phi}\\ \vdots\end{pmatrix}$. (23)

### v3

@@ Appendix A Application of Linear Optical Devices on Photonic States
When a single mode, carrying one or more photons, enters the interferometer, three ancillas in a vacuum state are added by the interferometric setup (see Figure 5). As mentioned above, the mode that enters the interferometer at time $t^{\prime}_{0}$, yields two modes at time $t_{0}$, and two modes at time $t_{1}$. These four output modes are: times $t_{0}$, $t_{1}$ at the ‘s’ (straight) arm of the interferometer, and times $t_{0}$, $t_{1}$ at the ‘d’ (down) arm of the interferometer. A basis state of this Fock-space can be written as $|{n_{d_{1}},n_{d_{0}},n_{s_{1}},n_{s_{0}}}\rangle^{\!\text{\tiny F}}$.
Refer to caption Pulse (1) is about to enter the interferometer. A vacuum ancilla (1’) is added at the input of the first beam splitter, $\mathrm{BS}_{1}$.
Refer to caption Pulses (1) and (1’) interfere in the first beam splitter ($\mathrm{BS}_{1}$) and yield a superposition of (2) and (3) in the short and long arms of the interferometer, respectively: $|{0}\rangle^{\!\text{\tiny F}}_{1^{\prime}}|{1}\rangle^{\!\text{\tiny F}}_{1}\stackrel{{\scriptstyle\mathrm{BS}_{1}}}{{\longrightarrow}}(|{0}\rangle^{\!\text{\tiny F}}_{3}|{1}\rangle^{\!\text{\tiny F}}_{2}+i|{1}\rangle^{\!\text{\tiny F}}_{3}|{0}\rangle^{\!\text{\tiny F}}_{2})/\sqrt{2}$. Pulse (2) is about to enter the second beam splitter ($\mathrm{BS}_{2}$), so a vacuum ancilla is added (2’).
Refer to caption Pulses (4) and (5) are created by pulses (2) and (2’): $\frac{1}{\sqrt{2}}|{1}\rangle^{\!\text{\tiny F}}_{2}|{0}\rangle^{\!\text{\tiny F}}_{2^{\prime}}\stackrel{{\scriptstyle\mathrm{BS}_{2}}}{{\longrightarrow}}(i|{0}\rangle^{\!\text{\tiny F}}_{5}|{1}\rangle^{\!\text{\tiny F}}_{4}+|{1}\rangle^{\!\text{\tiny F}}_{5}|{0}\rangle^{\!\text{\tiny F}}_{4})/2$. Pulse (3) is about to enter the second beam splitter, so a vacuum ancilla is added (3’).
Refer to caption Pulses (6) and (7) are created by the interference of (3) and (3’): $\frac{i}{\sqrt{2}}|{0}\rangle^{\!\text{\tiny F}}_{3^{\prime}}|{1}\rangle^{\!\text{\tiny F}}_{3}\stackrel{{\scriptstyle\mathrm{BS}_{2}}}{{\longrightarrow}}(i|{0}\rangle^{\!\text{\tiny F}}_{7}|{1}\rangle^{\!\text{\tiny F}}_{6}-|{1}\rangle^{\!\text{\tiny F}}_{7}|{0}\rangle^{\!\text{\tiny F}}_{6})/2$.
Figure 5. Evolution in time of a single photon pulse through the interferometer with $\phi=0$: $|{0{,}0{,}0{,}1}\rangle^{\!\text{\tiny F}}_{3^{\prime},2^{\prime},1^{\prime},1}\to\frac{1}{2}\left(|{0{,}0{,}0{,}1}\rangle^{\!\text{\tiny F}}-|{0{,}0{,}1{,}0}\rangle^{\!\text{\tiny F}}+i|{0{,}1{,}0{,}0}\rangle^{\!\text{\tiny F}}+i|{1{,}0{,}0{,}0}\rangle^{\!\text{\tiny F}}\right)_{6,4,7,5}$. The output state is denoted by modes $|{n_{d_{1}},n_{d_{0}},n_{s_{1}},n_{s_{0}}}\rangle^{\!\text{\tiny F}}$ that correspond to modes (6), (4), (7), and (5), respectively.
fully described inside the figure.
@@
(14) $|{0{,}0{,}0}\rangle^{\!\text{\tiny F}}|{1}\rangle^{\!\text{\tiny F}}_{t^{\prime}_{0}}\mapsto(|{0{,}0{,}0{,}1}\rangle^{\!\text{\tiny F}}-e^{i\phi}|{0{,}0{,}1{,}0}\rangle^{\!\text{\tiny F}}+i|{0{,}1{,}0{,}0}\rangle^{\!\text{\tiny F}}+ie^{i\phi}|{1{,}0{,}0{,}0}\rangle^{\!\text{\tiny F}})\thickspace/2\ .$
Note the three input vacuum ancillas that were added. Also note that a pulse sent at a different time (say, $t^{\prime}_{1}$ or $t^{\prime}_{-1}$) results in the same output state, with appropriate delays. That is, a single-photon pulse entering the interferometer at time $t^{\prime}_{k}$ results in the state $(|{0{,}0{,}0{,}1}\rangle^{\!\text{\tiny F}}-e^{i\phi}|{0{,}0{,}1{,}0}\rangle^{\!\text{\tiny F}}+i|{0{,}1{,}0{,}0}\rangle^{\!\text{\tiny F}}+ie^{i\phi}|{1{,}0{,}0{,}0}\rangle^{\!\text{\tiny F}})\thickspace/2$ in a Fock-space with basis states $|{n_{d_{k+1}},n_{d_{k}},n_{s_{k+1}},n_{s_{k}}}\rangle^{\!\text{\tiny F}}$.
Using the more standard notation of $\{|{t^{\prime}_{k}}\rangle\}$ input states and $\{|{s_{k}}\rangle,|{d_{k}}\rangle\}$ output states used in the rest of the paper, Eq. (14) is given by
(15) $|{t^{\prime}_{k}}\rangle\mapsto\left(|{s_{k}}\rangle+i|{d_{k}}\rangle-e^{i\phi}|{s_{k+1}}\rangle+ie^{i\phi}|{d_{k+1}}\rangle\right)/2\text{ .}$
@@
We are now ready to consider the setup of Figure 4 and two input modes, $t^{\prime}_{0}$ and $t^{\prime}_{1}$, that enter the interferometer one after the other, with exactly the same time difference $\Delta T$ as the interferometer’s arms. As a result of this precise timing, the two modes are transformed into a superposition of only six modes (instead of eight modes) at the outputs (see Figure 6). Four (vacuum state) ancillas are added during the process and the resulting six modes are $t_{0}$, $t_{1}$, $t_{2}$ at the ‘s’ arm and the ‘d’ arm of the interferometer. A basis state of this Fock-space is, therefore, $|{n_{d_{2}},n_{d_{1}},n_{d_{0}},n_{s_{2}},n_{s_{1}},n_{s_{0}}}\rangle^{\!\text{\tiny F}}$. If exactly one photon enters the interferometer, we can use Eq. (14) to obtain
@@
[…]
Recall that $|{0}\rangle=|{0{,}1}\rangle^{\!\text{\tiny F}}_{t^{\prime}_{1},t^{\prime}_{0}}$ and $|{1}\rangle=|{1{,}0}\rangle^{\!\text{\tiny F}}_{t^{\prime}_{1},t^{\prime}_{0}}$. It follows that an arbitrary qubit is transformed as
[…]
@@
Refer to caption A general single-photon qubit, $\alpha|{0{,}1}\rangle^{\!\text{\tiny F}}+\beta|{1{,}0}\rangle^{\!\text{\tiny F}}$, enters the interferometer (modes (2) and (1)). Bob adds a vacuum ancilla (1’) that interferes with mode (1) at the first beam splitter ($\mathrm{BS}_{1}$).
@@
Refer to caption Pulses (1) and (1’) interfere and yield pulses (3) and (4) in the short arm and the long arm, respectively: $\alpha|{0}\rangle^{\!\text{\tiny F}}_{1^{\prime}}|{1}\rangle^{\!\text{\tiny F}}_{1}\stackrel{{\scriptstyle\mathrm{BS}_{1}}}{{\longrightarrow}}\frac{\alpha}{\sqrt{2}}(|{0}\rangle^{\!\text{\tiny F}}_{4}|{1}\rangle^{\!\text{\tiny F}}_{3}+i|{1}\rangle^{\!\text{\tiny F}}_{4}|{0}\rangle^{\!\text{\tiny F}}_{3})$. Pulse (3) is about to enter $\mathrm{BS}_{2}$, so a vacuum ancilla (3’) is added. Pulse (2) is about to enter $\mathrm{BS}_{1}$, so a vacuum ancilla (2’) is added.
@@
Refer to caption Pulses (7) and (8) are created by the interference of (3) and (3’): $\frac{\alpha}{\sqrt{2}}|{1}\rangle^{\!\text{\tiny F}}_{3}|{0}\rangle^{\!\text{\tiny F}}_{3^{\prime}}\stackrel{{\scriptstyle\mathrm{BS}_{2}}}{{\longrightarrow}}\frac{\alpha}{2}(i|{0}\rangle^{\!\text{\tiny F}}_{8}|{1}\rangle^{\!\text{\tiny F}}_{7}+|{1}\rangle^{\!\text{\tiny F}}_{8}|{0}\rangle^{\!\text{\tiny F}}_{7})$. Pulses (5) and (6) are created by the interference of (2) and (2’) in $\mathrm{BS}_{1}$: $\beta|{0}\rangle^{\!\text{\tiny F}}_{2^{\prime}}|{1}\rangle^{\!\text{\tiny F}}_{2}\stackrel{{\scriptstyle\mathrm{BS}_{1}}}{{\longrightarrow}}\frac{\beta}{\sqrt{2}}(|{0}\rangle^{\!\text{\tiny F}}_{6}|{1}\rangle^{\!\text{\tiny F}}_{5}+i|{1}\rangle^{\!\text{\tiny F}}_{6}|{0}\rangle^{\!\text{\tiny F}}_{5})$.
@@
Refer to caption Pulses (9) and (10) are created by the interference of (4) and (5) in the second beam splitter: $\frac{i\alpha}{\sqrt{2}}|{0}\rangle^{\!\text{\tiny F}}_{5}|{1}\rangle^{\!\text{\tiny F}}_{4}+\frac{\beta}{\sqrt{2}}|{1}\rangle^{\!\text{\tiny F}}_{5}|{0}\rangle^{\!\text{\tiny F}}_{4}\stackrel{{\scriptstyle\mathrm{BS}_{2}}}{{\longrightarrow}}\frac{i(\alpha+\beta)}{2}|{0}\rangle^{\!\text{\tiny F}}_{10}|{1}\rangle^{\!\text{\tiny F}}_{9}+\frac{\beta-\alpha}{2}|{1}\rangle^{\!\text{\tiny F}}_{10}|{0}\rangle^{\!\text{\tiny F}}_{9}$ . Pulse (6) is about to enter $\mathrm{BS}_{2}$, so a vacuum ancilla is added (6’).
@@
Refer to caption Pulses (11) and (12) are created by the interference of (6) and (6’) in $\mathrm{BS}_{2}$: $\frac{i\beta}{\sqrt{2}}|{0}\rangle^{\!\text{\tiny F}}_{6^{\prime}}|{1}\rangle^{\!\text{\tiny F}}_{6}\stackrel{{\scriptstyle\mathrm{BS}_{2}}}{{\longrightarrow}}\frac{\beta}{2}(i|{0}\rangle^{\!\text{\tiny F}}_{12}|{1}\rangle^{\!\text{\tiny F}}_{11}-|{1}\rangle^{\!\text{\tiny F}}_{12}|{0}\rangle^{\!\text{\tiny F}}_{11})$.
@@
Figure 6. Evolution in time of two modes through the interferometer with $\phi=0$: $|{0{,}0{,}0{,}0}\rangle^{\!\text{\tiny F}}_{6^{\prime},3^{\prime},2^{\prime},1^{\prime}}\left(\alpha|{0}\rangle^{\!\text{\tiny F}}_{2}|{1}\rangle^{\!\text{\tiny F}}_{1}+\beta|{1}\rangle^{\!\text{\tiny F}}_{2}|{0}\rangle^{\!\text{\tiny F}}_{1}\right)\to\bigl(\frac{\alpha}{2}|{0{,}0{,}0{,}0{,}0{,}1}\rangle^{\!\text{\tiny F}}+\frac{\beta-\alpha}{2}|{0{,}0{,}0{,}0{,}1{,}0}\rangle^{\!\text{\tiny F}}-\frac{\beta}{2}|{0{,}0{,}0{,}1{,}0{,}0}\rangle^{\!\text{\tiny F}}+\frac{i\alpha}{2}|{0{,}0{,}1{,}0{,}0{,}0}\rangle^{\!\text{\tiny F}}+\frac{i(\alpha+\beta)}{2}|{0{,}1{,}0{,}0{,}0{,}0}\rangle^{\!\text{\tiny F}}+\frac{i\beta}{2}|{1{,}0{,}0{,}0{,}0{,}0}\rangle^{\!\text{\tiny F}}\bigr)_{11,9,7,12,10,8}$. The output state is denoted by modes $|{n_{d_{2}},n_{d_{1}},n_{d_{0}},n_{s_{2}},n_{s_{1}},n_{s_{0}}}\rangle^{\!\text{\tiny F}}$.
fully described inside the figure.
@@
While a pulse can arrive at the interferometer at any instance of time (which is continuous), we know that pulses only interfere with each other if their time difference is equal to the time difference of the interferometer’s two paths. Thus, we define some instance of time as $t=0$, and consider all times $\{t_{n}\}_{n}=\{n\cdot\Delta T\ \;|\;n\in\mathbb{Z}\}$, where $\Delta T$ is the interferometer’s time difference.
@@
(18) $\operatorname{span}\{|{a_{n}}\rangle,|{b_{n}}\rangle\;|\;n\in\mathbb{Z}\},$
@@
(19) $\operatorname{span}\{|{s_{n}}\rangle,|{d_{n}}\rangle\;|\;n\in\mathbb{Z}\},$
while we have already shown that the evolution for $|{a_{n}}\rangle$ is given by:
(20) $U|{a_{n}}\rangle=\left(|{s_{n}}\rangle+i|{d_{n}}\rangle-e^{i\phi}|{s_{n+1}}\rangle+ie^{i\phi}|{d_{n+1}}\rangle\right)/\thickspace 2.$
@@
(22) $U=\frac{1}{2}\begin{pmatrix}1&i&0&0&\cdots\\ i&-1&0&0\\ -e^{i\phi}&ie^{i\phi}&1&i\\ ie^{i\phi}&e^{i\phi}&i&-1\\ 0&0&-e^{i\phi}&ie^{i\phi}\\ 0&0&ie^{i\phi}&e^{i\phi}\\ \vdots&&&\end{pmatrix}$.
@@
(23) $U^{-1}=U^{\dagger}=\frac{1}{2}\begin{pmatrix}1&-i&-e^{-i\phi}&-ie^{-i\phi}&0&0&\cdots\\ -i&-1&-ie^{-i\phi}&e^{-i\phi}&0&0\\ 0&0&1&-i&-e^{-i\phi}&-ie^{-i\phi}\\ 0&0&-i&-1&-ie^{-i\phi}&e^{-i\phi}\\ \vdots\end{pmatrix}$.